# A waveform and time digitization mainboard prototype for the hybrid digital optical module of TRIDENT neutrino experiment

### v2

@@ 4 Summary
In summary, we presented a first fully custom-designed waveform and time digitization mainboard prototype for the hDOM of the planned TRIDENT neutrino experiment in the West Pacific Ocean. The prototype system demonstrates three key capabilities: (1) simultaneous waveform digitization of up to 32 PMT channels at 125 MS/s sampling rate; (2) single photoelectron detection with clear separation from pedestal noise, while maintaining good linearity up to 240 PEs, 60% higher than that of the mainboard for IceCube Upgrade; and (3) sub-nanosecond timing resolution for 32 PMT and 24 SiPM pulse leading edges, achieved through FPGA-based time-to-digital converter implementation. These performance characteristics make this mainboard suitable for the first hDOM prototype string sea trial planned in Fall 2025.
@@ 5 Acknowledgement
This work has received support from the Ministry of Science and Technology of China under Grant No.2022YFA1605500, the Office of Science and Technology of the Shanghai Municipal Government under Grant No.22JC1410100, and support from the Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education. We thank Xin Xiang to provide the picosecond laser source for the linearity evaluation. We thank Hualin Mei to review and improve this paper.

### v3

@@ 4 Summary
In summary, we presented a first fully custom-designed waveform and time digitization mainboard prototype for the hDOM of the planned TRIDENT neutrino experiment in the West Pacific Ocean. The prototype system demonstrates three key capabilities: (1) simultaneous waveform digitization of up to 32 PMT channels at 125 MS/s sampling rate; (2) single photoelectron detection with clear separation from pedestal noise, while maintaining good linearity up to 240 PEs; and (3) sub-nanosecond timing resolution for 32 PMT and 24 SiPM pulse leading edges, achieved through FPGA-based time-to-digital converter implementation. These performance characteristics make this mainboard suitable for the first hDOM prototype string sea trial planned in Fall 2025.
@@ 5 Acknowledgement
This work has received support from the Ministry of Science and Technology of China under Grant No.2022YFA1605500, the Office of Science and Technology of the Shanghai Municipal Government under Grant No.22JC1410100, and support from the Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education. We thank Xin Xiang to provide the picosecond laser source for the linearity evaluation. We thank Hualin Mei to review and improve this paper. On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data sets generated during the current study are available from the corresponding author on reasonable request.